OpenStudy (kellycahill):

what is the volume of the pyramid? base is 16ft, length is 16ft, height is 17ft A) 1,280ft^3 B)34,816ft^3 C)8,704ft^3 D)4,624ft^3

OpenStudy (photon336):

do you have the formula for a pyramid volume?

OpenStudy (astrophysics):

Assuming it's right rectangular I'd say \[V = \frac{ lwh }{ 3 }\]

OpenStudy (astrophysics):

So plug and chug at this point

OpenStudy (kellycahill):

I don't have the formula

OpenStudy (kellycahill):

the answer I got isn't an option

OpenStudy (astrophysics):

Ok I see you need the slant height, meaning we need surface area

OpenStudy (kellycahill):

I am guessing so I know nothing on this hahahaha

jimthompson5910 (jim_thompson5910):

it says `what is the volume of the pyramid?` so I'm not sure where you're getting surface area from. It could be a typo though.

OpenStudy (astrophysics):

Yeah sorry, I meant to get slant height we can do the following, \[S = \sqrt{r^2+h^2}\] where S is the slant height

OpenStudy (astrophysics):

Lately I don't know if I'm just reading these problems wrong or it's just worded weirdly, so we have the slant height, we want the "altitude height" then you can solve for h using the formula above, getting \[h = \sqrt{S^2-r^2}\] and yeah ok this makes more sense, we have \[h = \sqrt{17^2-8^2} = 15\] now we have our volume \[V = \frac{ hlw }{ 3 }\] we just solved for h, now plug in the l and w, haha.